英語 (English)

発信型英語 (Communicative English) (工 ((化 B ー 2)2 年))

スタージドナルド · 講師/ 大学院ソシオ・アーツ・アンド・サイエンス研究部 2 単位後期火 1 ・ 2

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 The primary aim of this course is to improve the student’s ability

to communicate in English in a variety of formal and informal situations. 【授業の概要】 There will be lectures, video presentations, and reading and writing

exercises, but the main focus will be on speaking and listening exercises. Class presentations will be an important component, and students will be expected to speak in class and to prepare written summaries of oral presentations given by other students.

【キーワード】 Communicative English 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】 Students who complete this course will have experience using English

to describe orally and in writing experiences and opinions of themselves and others. Oral presentation exercises should make students more comfortable in public speaking. In addition, students should have increased confidence in their foreign language ability because they have followed a course given by a foreign instructor entirely in English.

【授業の計画】 1. Course Outline

2. Self-Introduction

3. Discussion on Selected Topic(s)/Video

4. Class Presentations/Summarizing

5. Class Presentations/Summarizing

6. Discussions on Selected Topic(s)/Video

7. Discussions on Selected Topic(s)/Video

8. Mid Term Exam

9. Class Presentations/Summarizing

10. Class Presentations/Summarizing

11. Discussion on Selected Topic(s)/Video

12. Discussion on Selected Topic(s)/Video

13. Discussion on Selected Topic(s)/Video 14. Class Presentations/Summarizing/Course Evaluation

15. Class Presentations/Summarizing/Course Evaluation

16. [Note: This schedule is tentative and may be modified due to class size or other factors.]

【教科書】 There will be teacher handouts and text materials created by students. Students may also be expected to purchase copies, of newspapers for class activities.

【参考書等】 [参考資料]

【成績評価の方法】 Class presentations and written summaries of presentations, oral/written examinations and participation will be used for evaluation. Tentative evaluation scheme: Presentations=30% (3 X 10%), Examinations=20% (2 X 10%), Participation=50%.

【再試験の有無】 There is no make-up examination for this course.

【受講者へのメッセージ】 Class structure depends on the number of students enrolled. It may be possible to divide the class into smaller groups to facilitate conversation activities.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221221 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】

⇒ スタージ (総合科学部 1 号館 2303, 088-656-7134, dws@ias.tokushima-u.a c.jp) ���� (オフィスアワー: 水曜日 16:30-17:30 または応相談)

English

Communicative English (工 ((化 B ー 2)2 年))

Donald Sturge · Associate Professor / Institute of Socio-Arts and Sciences 2 units 後期火 1 ・ 2

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ The primary aim of this course is to improve the student’s ability to

communicate in English in a variety of formal and informal situations. Outline⟩ There will be lectures, video presentations, and reading and writing

exercises, but the main focus will be on speaking and listening exercises. Class presentations will be an important component, and students will be expected to speak in class and to prepare written summaries of oral presentations given by other students.

�������⟩ Communicative English Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ Students who complete this course will have experience using English to

describe orally and in writing experiences and opinions of themselves and others. Oral presentation exercises should make students more comfortable in public speaking. In addition, students should have increased confidence in their foreign language ability because they have followed a course given by a foreign instructor entirely in English.

Schedule⟩

1. Course Outline

2. Self-Introduction

3. Discussion on Selected Topic(s)/Video

4. Class Presentations/Summarizing

5. Class Presentations/Summarizing

6. Discussions on Selected Topic(s)/Video

7. Discussions on Selected Topic(s)/Video

8. Mid Term Exam

9. Class Presentations/Summarizing

10. Class Presentations/Summarizing

11. Discussion on Selected Topic(s)/Video

12. Discussion on Selected Topic(s)/Video

13. Discussion on Selected Topic(s)/Video 14. Class Presentations/Summarizing/Course Evaluation

15. Class Presentations/Summarizing/Course Evaluation

16. [Note: This schedule is tentative and may be modified due to class size or other factors.]

Textbook⟩ There will be teacher handouts and text materials created by students. Students may also be expected to purchase copies, of newspapers for class activities.

Reference⟩ [参考資料]

Evaluation Criteria⟩ Class presentations and written summaries of presentations, oral/written examinations and participation will be used for evaluation. Tentative evaluation scheme: Presentations=30% (3 X 10%), Examinations=20% (2 X 10%), Participation=50%.

Re-evaluation⟩ There is no make-up examination for this course.

Message⟩ Class structure depends on the number of students enrolled. It may be possible to divide the class into smaller groups to facilitate conversation activities.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221221 Contact (Office-Hour, Room, E-mail)⟩

⇒ Sturge (Bldg. #1 2303, +81-88-656-7134, dws@ias.tokushima-u.ac.jp) ���� (Office Hour: Wednesday from 16:30-17:30 or by appointment)